# Linear average-case complexity of algorithmic problems in groups

### v5

@@ Abstract.
Research of the first author was partially supported by the NSF award DMS-1901976. Research of the second author was partially supported by the RFBR Grant 20-01-00465
@@ 4. The word problem in free products of groups
We may assume that at least one of the groups $A$ and $B$ has more than 2 elements. Let the group $A$ be generated by $a_{1},\ldots,a_{s}$ and the group $B$ by $b_{1},\ldots,b_{t}$.
@@
Suppose that after $n$ stages of the algorithm $\mathcal{A}$, the prefix $v=w(n)$ of length $n$ of the word $w$ is already in the form $w_{1}w_{2}\cdots w_{i}$ as above. Let $x$ be the $(n+1)$th letter of the word $w$. Then the $(n+1)$th stage of the algorithm $\mathcal{A}$ proceeds as follows. If $x$ is not an element of the alphabet that $w_{i}$ is a word in, then we check if $x=1$. If it is, then the form for the prefix $w(n+1)$ does not change. If $x\neq 1$, then $w_{i+1}=x$.
@@
(b) The number of words of length $r$ equal to 1 in the group $G=A\ast B$ is at most $(2m)^{r}\cdot\exp(-c_{2}\cdot r)$, where $m=s+t$ is the number of generators of $G$ and $c_{2}$ is a positive constant.
@@
There is a positive constant $c$ such that the number of simple words $w$ of length $k$ in the $A$-alphabet is at most $2(2m)^{k}\exp(-c\cdot p)$, where $p=\lfloor\sqrt{k}\rfloor$, the floor function of $\sqrt{k}$.
@@
By the definition of $p$, it is possible to split the word $w$ into subwords $w_{1},\ldots,w_{l}$ of fixed lengths $p_{i}\geq p$, where $l\geq p$. It is sufficient to show that, for large enough $k$ and $p$, there are no more than $(2m)^{p_{i}}\lambda$ ways to select each $w_{i}$, for some constant $\lambda<1$. Indeed, in that case the number of different words $w$ will be no greater than $(2m)^{k}\lambda^{p}$, as desired.
@@
Let $v_{i}$ and $v_{i}^{\prime}$ be two words like that. If the last letter of the geodesic form $g_{i}$ of $v_{i}$ and the first letter of the geodesic form $g_{i}^{\prime}$ of $v_{i}^{\prime}$ are from different alphabets, then, if $g_{i}$ and $g_{i}^{\prime}$ are of length $\geq c_{1}s_{i}$, the length of the geodesic word $g_{i}g_{i}^{\prime}$ has length $>c_{1}s_{i}$ and has a letter $b$ from the $B$-alphabet surrounded by letters of the $A$-alphabet.
This letter $b$ will still be there after any reduction of the word $u_{i}=v_{i}v_{i}^{\prime}$ to the standard form, and moreover the letter $b$ will be at a distance $\geq c_{1}s_{i}-1$ from the beginning as well as from the end of $u_{i}$.
@@
For any words $w_{i}$ obtained as above, this letter $b$ will not cancel out upon any reduction of $w=w_{1}\cdots w_{d}$ to the standard form. Indeed, for $b$ to cancel out, there should be a subword of $w$ equal to 1 in $G$ and having length $\geq c_{1}p$, but there are no subwords like that. Thus, we get a contradiction with the simplicity of the word $w$.
@@
Let now $M(k)$ be the set of all group words in the given alphabet with $m$ letters such that the number of words of length $k$ in $M(k)$ is at most $\exp(-cp)(2m)^{k}$ for some $c>0$ (provided by Lemma 7), where $p=\lfloor\sqrt{k}\rfloor$.
@@
if $C$ is selected large enough compared to $d$. This completes the proof of part (b) of the lemma.
@@
2. Let $V_{n}^{(2)}$ denote the set of words $w$ of length $n$ that are not in $V_{n}^{(1)}$ and for any integer $k$ in the interval $(C,C(\log n)^{2})$, $w$ contains less than $[n/k^{d+3}]$ simple subwords of length $k$.
Then by Lemma 4, for any integer $k$ in the interval $(C,C(\log n)^{2})$, the reduced form of the word $w$ will have less than $[n/k^{d+3}]$ syllables of length $k$ produced by the algorithm $\mathcal{A}$ while working on subwords of length $k<C(\log n)^{2}$. Therefore, producing these syllables takes time $\leq K(k^{d+1}[n/k^{d+3}]<2Kn/k^{2}$. Summing up over all relevant values of $k$, we see that producing all syllables of length $\geq C$ takes time $O(n)$.
@@
3. All the remaining words of length $n$ form a set that we denote by $V_{n}^{(3)}$. This set is the union, over all $k$, $C<k<C(\log n)^{2}$, of the subsets $M(n,k,l)$, where $l=[n/k^{d+3}]$.
@@
Summing up, we see that the average-case time complexity of the algorithm $\mathcal{A}$ is $O(n)$, and this completes the proof.
$\Box$
@@ 5. The subgroup membership problem in free products of groups
(b) Suppose a subgroup $H\subset G$ has trivial intersection with any subgroup of $G$ conjugate (in $G$) to $A$ or $B$. Then the (worst-case) time complexity of the subgroup membership problem in $H$ is not higher than the time complexity of the word problem in $G$.
@@ 6. The identity problem
and this completes the proof of the lemma. ∎
We now continue with the proof of Theorem 9. In reference to Lemma 9, let $k<c\log_{4}n$ for some constant $c,~{}0<c<0.5$. Then
@@
Now recall that there are $2^{k-1}$ different letter/inverse letter pairs of complexity $k$. Therefore, with $k<c\log_{4}n$, for the set $Z_{n}$ of all words of complexity $n$ that do not contain at least one letter $x_{i}^{\pm 1}$ of complexity $k$ we have
@@
for sufficiently large $n$. Therefore, $\frac{|W_{n}-Z_{n}|}{|W_{n}|}>1-\exp(-n^{c})$.
@@ 7. Open problems
    We note that the converse is not true, see the comment to the Problem 1 above. We also note that if the sums in the formula (1) are bounded by a linear function of $n$, then for any superlinear function $f(n)$ the densities of subsets of $G$ where $T(w)>f(n)$ (for $w\in W_{n}$) are approaching 0 when $n$ goes to infinity. This means that if the average-case time complexity of the word problem in $G$ is linear, then the generic-case time complexity of the word problem in $G$ can be bounded by any superlinear function.

### v6

@@ Abstract.
Research of the first author was partially supported by the NSF award DMS-1901976.
@@ 4. The word problem in free products of groups
We may assume that at least one of the groups $A$ and $B$ has more than 2 elements. Let the group $A$ be generated by nontrivial $a_{1},\ldots,a_{s}$ and the group $B$ by nontrivial $b_{1},\ldots,b_{t}$.
@@
Suppose that after $n$ stages of the algorithm $\mathcal{A}$, the prefix $v=w(n)$ of length $n$ of the word $w$ is already in the form $w_{1}w_{2}\cdots w_{i}$ as above. Let $x$ be the $(n+1)$th letter of the word $w$. Then the $(n+1)$th stage of the algorithm $\mathcal{A}$ proceeds as follows. If $x$ is not an element of the alphabet that $w_{i}$ is a word in, then $w_{i+1}=x$.
@@
(b) The number of words of length $r$ equal to 1 in the group $G=A\ast B$ is at most $(2m)^{r}\cdot\exp(-c_{2}\cdot r)$, where $m=s+t$ is the number of generators of $G$ and $c_{2}$ is a positive constant independent of $r$.
@@
There is a positive constant $c$ such that the number of simple words $w$ of any length $k$ of type $A$ is at most $(2m)^{k}\exp(-c\cdot p)$, where $p=\lfloor\sqrt{k}\rfloor$, the floor function of $\sqrt{k}$.
@@
By the definition of $p$, it is possible to split the word $w$ into subwords $w_{1},\ldots,w_{l}$ of fixed lengths $p_{i}\geq p$, where $l\geq p$. It is sufficient to show that, for large enough $k$ and $p$, there are no more than $(2m)^{p_{i}}\lambda$ ways to select each $w_{i}$, for some constant $\lambda<1$. Indeed, in that case the number of different words $w$ will be not greater than $(2m)^{k}\lambda^{p}$, as desired.
@@
Let $v_{i}$ and $v_{i}^{\prime}$ be two words like that. If the last letter of the geodesic form $g_{i}$ of $v_{i}$ and the first letter of the geodesic form $g_{i}^{\prime}$ of $v_{i}^{\prime}$ are from different alphabets, then, if $g_{i}$ and $g_{i}^{\prime}$ are of length $\geq c_{1}s_{i}$, the length of the geodesic word $g_{i}g_{i}^{\prime}$ has length $>c_{1}s_{i}$ and has a letter $b$ from the $B$-alphabet at the concatenation of $g_{i}$ and $g_{i}^{\prime}$.
A letter from the $B$-alphabet (we keep the same notation $b$ for it) will still be there after any reduction of the word $u_{i}=v_{i}v_{i}^{\prime}$ to the standard form, and moreover the letter $b$ will be at a distance $\geq c_{1}s_{i}-1$ from the beginning as well as from the end of $u_{i}$.
@@
For any words $w_{i}$ obtained as above, this letter $b$ will not cancel out upon any reduction of $w=w_{1}\cdots w_{d}$ to the standard form. Indeed, for $b$ to cancel out, there should be a subword of $w$ equal to 1 in $G$ and having length $\geq c_{1}p$, but there are no subwords like that. Thus, we get a contradiction with the simplicity of type $A$ of the word $w$.
@@
Let now $M(k)$ be the set of all simple group words of length $k$ in the given alphabet with $m$ letters. By Lemma 7, the cardinality of $M(k)$ is at most $\exp(-cp)(2m)^{k}$ for some $c>0$, where $p=\lfloor\sqrt{k}\rfloor$.
@@
if $C$ is selected large enough compared to both $c$ and $d$. This completes the proof of part (b) of the lemma.
@@
2. Let $V_{n}^{(2)}$ denote the set of words $w$ of length $n$ that are not in $V_{n}^{(1)}$ and for any integer $k<C(\log n)^{2}$, $w$ contains less than $[n/k^{d+3}]$ simple subwords of length $k$.
Then by Lemma 4, for any integer $k$ in the interval $(C,C(\log n)^{2})$, the reduced form of the word $w$ will have less than $[n/k^{d+3}]$ syllables of length $k$ produced by the algorithm $\mathcal{A}$ while working on subwords of length $k<C(\log n)^{2}$. Therefore, producing these syllables takes time $\leq K(k^{d+1}[n/k^{d+3}])<2Kn/k^{2}$. Summing up over all relevant values of $k$, we see that producing all syllables of length $\geq C$ takes time $O(n)$.
@@
3. All the remaining words of length $n$ form a set that we denote by $V_{n}^{(3)}$. This set is the union, over all $k$, $C<k<C(\log n)^{2}$, of the subsets $M(n,k,l)$, where $l=\lfloor n/k^{d+3}\rfloor$.
@@
Summing up, we see that the average-case time complexity of the algorithm $\mathcal{A}$ is $O(n)$, and this completes the proof. $\Box$
@@ 5. The subgroup membership problem in free products of groups
(b) Suppose a subgroup $H\subset G$ has trivial intersection with any subgroup of $G$ conjugate (in $G$) to $A$ or $B$. Then the (worst-case) time complexity of the subgroup membership problem in $H$ is not higher than the time complexity of putting words into free product normal form in $G$.
@@ 6. The identity problem
and this completes the proof of the lemma. (For the latter inequality we use the fact that $(1-x)^{n}>1-nx$ if $x\in(0,1)$ and the integer $n$ is greater than 1.) ∎
We now continue with the proof of Theorem 9. In reference to Lemma 9, let $k\leq c\log_{4}n$ for some constant $c,~{}0<c<0.5$. Then
@@
Now recall that there are $2^{k-1}$ different letter/inverse letter pairs of complexity $k$. Therefore, with $k\leq c\log_{4}n$, for the set $Z_{n}$ of all words $w$ of complexity $n$ such that there is at least one letter $x_{i}^{\pm 1}$ of complexity $k$ that does not occur in $w$, we have
@@
for sufficiently large $n$. (The last inequality follows from the fact that $n^{c/2}\exp(n^{c}-n^{(2-c)/3})$ converges to 0 when $n\to\infty$ since $c<0.5.$) Therefore, $\frac{|W_{n}-Z_{n}|}{|W_{n}|}>1-\exp(-n^{c})$.
@@ 7. Open problems
    We note that if the sums in the formula (1) are bounded by a linear function of $n$, then for any superlinear function $f(n)$ the densities of subsets of $G$ where $T(w)>f(n)$ (for $w\in W_{n}$) are approaching 0 when $n$ goes to infinity. This means that if the average-case time complexity of the word problem in $G$ is linear, then the generic-case time complexity of the word problem in $G$ can be bounded by any superlinear function.